y"(t) + 4y'(t) + 4y(t) = 4e-2t, y(0) = -1, y'(0) = 4


1
Expert's answer
2021-07-14T14:44:38-0400

Characteristic equation corresponding to the given homogeneous differential equation

"y''(t)+4y'(t)+4y=0" is "r^2+4r+4=0"

"\Rightarrow (r+2)^2=0\Rightarrow r=-2,-2"

General solution to the homogeneous differential equation is

"y_{h}=c_{1}e^{-2t}+c_2te^{-2t}"

Use Method of variation of parameters to find the particular solution

Particular solution to the given differential equation is

"y_{p}=Au+Bv,u=e^{-2t},v={te^{-2t}}" where

"A=\int \frac{-vg(t)}{uv'-vu'}dt,B=\int \frac{ug(t)}{uv'-vu'}dt,g(t)=4e^{-2t}"

Find "uv'-vu'"

Now

"uv'-vu'=(e^{-2t})\left [ e^{-2t}-2te^{-2t} \right ]-(te^{-2t})(-2e^{-2t})=e^{-4t}-2te^{-4t}+2te^{-4t}=e^{-4t}"

"A=\int \frac{-e^{-2t}(4e^{-2t})}{e^{-4t}}dt=-4\int dt=-4t"

"B=\int \frac{te^{-2t}(4e^{-2t})}{e^{-4t}}dt=4\int(t) dt=\frac{t^2}{2}"

Using the above, we get

"y_{p}=(-4t)(e^{-2t})+(\frac{t^2}{2})({te^{-2t}})=-4te^{-2t}+\frac{t^3}{2}e^{-2t}"

General solution to the given differential equation is

"y=y_{h}+y_{p}=c_{1}e^{-2t}+c_{2}te^{-2t}-4te^{-2t}+\frac{t^3}{2}e^{-2t}"

Find the value of "c_{1}" and "c_{2}" using the given initial conditions "y(0)=-1,y'(0)=4"

"y=c_{1}e^{-2t}+c_{2}te^{-2t}-4te^{-2t}+\frac{t^3}{2}e^{-2t}"

"\Rightarrow y'(t)=-2c_{1}e^{-2t}+c_{2}(e^{-2t}-2te^{-2t})-4e^{-2t}+8te^{-4t}+\frac{3t^2}{2}e^{-2t}-t^3e^{-2t}"

"y(0)=-1\Rightarrow c_{1}+0=-1\Rightarrow c_{1}=-1"

"y'(0)=4\Rightarrow -2c_{1}+c_{2}(1-0)-4(1)+8(0)+0-0=4"

"\Rightarrow -c_{1}+c_{2}=0\Rightarrow c_{2}=-1"

Therefore, solution to the given differential equation is

"y=-e^{-2t}-te^{-2t}-4te^{-2t}+\frac{t^3}{2}e^{-2t}"